yomomalikebracobama:

Solve for x: the quantity of x plus 4 over 3 = 2. (1 point) a x = −2 b x = 2 c x = 2 over 3 d x = negative 10 over 3

QuestionCoveBot:

Equation looks like: \[x+\frac{ 4 }{ 3 }=2\]

yomomalikebracobama:

solve for x :x+4/3 =2

yomomalikebracobama:

thats what it looks like

yomomalikebracobama:

so how would i find x?

yomomalikebracobama:

mr math robot i shall call you

QuestionCoveBot:

So first we have to subtract... \[\frac{ 4 }{ 3 }\]... on both sides. It would then look like; \[x+\frac{ 4 }{ 3 }-\frac{ 4 }{ 3 }=2-\frac{ 4 }{ 3 }\] And please, call me QCB.

QuestionCoveBot:

So what is 2-4/3 ?

yomomalikebracobama:

idk i dum

yomomalikebracobama:

ik we divide right?

QuestionCoveBot:

No we subract 2 drom 4 over 3. So what is... \[2-\frac{ 4 }{ 3 }\]

yomomalikebracobama:

what is drom

QuestionCoveBot:

*from sorry

yomomalikebracobama:

so 2/2?

QuestionCoveBot:

No. It would be 2/3 in fraction form, but .666 in decimal form.

yomomalikebracobama:

oooo

yomomalikebracobama:

i see me made mistake

QuestionCoveBot:

But we need fraction form. So after subtracting, we end up with... \[x=\frac{ 2 }{ 3 }\]